# KAN KAN Buff Signed Graph Neural Networks?

### v1

@@ Abstract
Graph Representation Learning aims to create embeddings for nodes and edges, capturing their features and interconnections. Graph Neural Networks (GNNs) have excelled in this task, leveraging neural networks to model complex graph relationships. Recently, the Kolmogorov-Arnold Neural Network (KAN) emerged as an alternative to Multi-Layer Perceptron (MLP), showing improved accuracy and interpretability with fewer parameters. While KANs have been integrated into unsigned GNNs, their application in signed GNNs remains unexplored. This paper integrates KAN into Signed Graph Convolutional Networks (SGCNs) to evaluate its performance on signed graphs where edges have positive or negative signs. We empirically assess KAN-enhanced SGCNs (KASGCN) on downstream tasks such as signed community detection and link sign prediction to enhance the embedding quality in signed networks. Considering the variability in the results indicated by the relatively large standard deviation, KASGCN demonstrates competitive performance with, or similar to, the vanilla SGCN in the evaluated downstream tasks, and its effectiveness is context-dependent (signed graph and parameters…etc.).
@@
1 Introduction
Graph Representation Learning is the process of fitting embeddings of nodes/edges based on the features and connections of the graph. These embeddings capture the local and global features of the nodes and edges in the original graph and are used in many downstream tasks such as node classification and link prediction [Xu, 2021]. One of the most prominent methods for learning graph embeddings is Graph Neural Networks (GNNs). They have shown remarkable effectiveness in capturing graphs’ complex relationships and structures [Fan et al., 2019, Bessadok et al., 2023]. GNNs utilize neural networks to gather and process information from a node’s surrounding nodes, gradually creating more sophisticated embeddings through successive layers.
Recently, the Kolmogorov-Arnold Neural Network (KAN) emerged as an alternative to Multi-Layer Perceptron (MLP) [Liu et al., 2024] and has demonstrated promising results in various benchmarks. Inspired by the Kolmogorov-Arnold representation theorem, KANs have learnable activation functions on edges, replacing fixed activation functions on nodes. This method has improved accuracy and interpretability, often surpassing MLPs with fewer parameters and more compact network architectures. Consequently, there have been multiple attempts to explore, validate, and integrate KANs in a variety of applications and architectures, which encompasses graph neural networks [De Carlo et al., 2024, Kiamari et al., 2024, Bresson et al., 2024] in a bid to achieve better performance and improve the quality of the embeddings. However, all the related work focuses on employing KANs in unsigned GNNs, and there is no research done regarding the effect of KANs on signed GNNs, which are designed to learn embeddings for signed graphs where edges can have positive or negative signs indicating different types of interactions.
@@ 1 Introduction
In this paper, we adopt the most basic variant of signed GNNs proposed by Derr et al. [Derr et al., 2018], which is a Signed Graph Convolutional Neural Network (SGCN) for simplicity and as a baseline. This variant of GNN was the first to adapt the Graph Convolutional Neural Network for signed networks based on the balance theory. It is a pivotal concept that explains the evolution of attitudes within networks. Established by Heider [Abelson and Rosenberg, 1958] and subsequently formalized mathematically by Harary, who introduced $k$-way balancing [Cartwright and Harary, 1956, Harary and Cartwright, 1968], balance theory has applications in predicting edge sentiment, content and product recommendations, and anomaly detection in multiple domains [Derr et al., 2020, Garimella et al., 2021, Interian et al., 2022, Amelkin and Singh, 2019]. A signed network is considered strongly balanced if every fundamental cycle is positive (the number of negative edges for every cycle is even). We are the first to integrate KAN (with its variants) into signed graph neural networks and empirically evaluate its impact on performance on downstream tasks (signed clustering and link sign prediction) for signed networks. In addition, this study aims to explore the potential of KANs in increasing the expressive power of the embeddings generated by signed graph neural networks.
@@
The paper is organized as follows: Section 2 discusses the related work regarding KAN integration into several architectures and their uses. In Section 3, we present the methodology of our work. In Section 4, we empirically evaluate the KAN-integrated signed GNN and its variants in several tasks, including efficiency examination and embedding comparison. We structure this section by presenting the research question, hypothesis, results, and observations. In Section 5, we summarize our findings.
2 Related Work
@@
Cheon et al. [Cheon, 2024] introduced a pioneering method that integrates Kolmogorov-Arnold Networks (KANs) with various pre-trained Convolutional Neural Network models, such as VGG16 and MobileNetV2, for Remo, the sensing scene classification using the Eurostat dataset. This approach achieved high accuracy with fewer epochs and parameters than traditional Multi-Layer Perceptrons without KAN integration. Bresson et al. proposed to integrate the KAN layer into unsigned graph neural networks (GNNs) and evaluated its performance in the node and graph classification tasks [Bresson et al., 2024]. Li et al. utilized Fourier series-based KAN’s power to optimize GNNs, and the evaluation took place on datasets for molecular property prediction [Li et al., 2024]. KANs were also synergized with science by using a pykan package that contains many functionalities, like compiling symbolic formulas into KANs. Based on these tools, KANs’ ability to detect many physical and constitutive laws is demonstrated [Liu et al., 2024]. In addition, Dong et al. evaluated the performance of KANs on time series data and found that KANs can match or even surpass the performance of MLP across 128 different time series datasets [Dong et al., 2024]. KANs have been explored for data representation using autoencoders and compared against traditional Convolutional Neural Networks (CNNs) on datasets such as MNIST, SVHN, and CIFAR-10. And it is shown that KAN-based autoencoders deliver competitive reconstruction accuracy [Moradi et al., 2024]. Lu et al. assessed using KANs for fraud detection and concluded that their performance varies depending on the context [Lu and Zhan, 2024]. Finally, the study in [Anonymous, 2024] presented a proof of concept for employing KANs in graphs to predict molecules’ binding affinity to protein targets. The model in this study did not attain state-of-the-art performance, but it emphasized its promising method in computational drug discovery.
3 Methodology
3.1 Kolmogorov-Arnold Networks Background
@@
According to the Kolmogorov-Arnold representation theorem [Liu et al., 2024], any multivariate function can be formulated as a combination of continuous univariate functions. It can be represented as:
@@
where $\alpha$ and $\phi$ are the univariate functions and $d$ is the dimension of the input. A univariate and non-linear spline function is applied for each input dimension. The output of each is aggregated. The main hyperparameters for the spline function are the spline order (degree of B-splines) and the grid size (number of intervals to approximate the real function). In the original paper [Liu et al., 2024], KANs demonstrated greater expressiveness than standard multi-layer perceptrons (MLPs), and KANs surpassed MLPs with significantly fewer parameters. In MLPs, weight adjustments occur globally, which can lead to forgetting previously learned data when new data is introduced. Conversely, in KANs, modifying spline control points impacts only local regions (locality property), thereby preserving previous training. $\phi(x)$ is formulated as:
$\phi(x)=w_{b}b(x)+w_{s}\ \text{spline}(x).$ (2)
$b(x)=\text{silu}(x)=\frac{x}{1+e}$ (3)
$\text{spline}(x)=\sum_{i}c_{i}B_{i}(x)$ (4)
where $b(x)$ is the SiLU activation function, $\text{spline}(x)$ is a linear combination of B-splines, $c_{i}$ is a trainable parameter and $w_{b}$ and $w_{s}$ are trainable and control the magnitude of the activation function. $w_{s}=1$ and $w_{b}$ is initialized using Xavier initialization.
3.2 Adaptation of Signed Graph Convolutional Network to Kolmogorov-Arnold Signed Graph Convolutional Network
Derr. et al. [Derr et al., 2018] employ balance theory to aggregate features from a node’s neighbors appropriately. Each node in the signed network has a dual representation (positive and negative representation). During message passing, to update the positive representation of a node, the positive embeddings are aggregated from the positive neighbors, and the negative embeddings are aggregated from the negative neighbors to be concatenated with each other along its positive representation. On the other hand, for updating the negative representation of a node, the negative embeddings are aggregated from the positive neighbors, and the positive embeddings are aggregated from the negative neighbors to be concatenated with each other along its negative representation. Finally, the positive and negative embeddings are concatenated to obtain the final embedding. The signed GCN formulation is as follows:
@@
$h_{i}^{B(1)}=\sigma(W^{B(1)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (5)
$h_{i}^{U(1)}=\sigma(W^{U(1)}[\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (6)
@@
$h_{i}^{B(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}])$ (7)
$h_{i}^{U(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}])$ (8)
where $l$ is the layer number, $h_{i}^{B(l)}$ and $h_{i}^{U(l)}$ are the positive and negative representations of node $i$ in layer $l$ respectively, $\sigma$ is an activation function, $W^{B(l)}$ and $W^{U(l)}$ are the weight matrices for updating the positive and negative representations respectively, $N_{i}^{+}$ and $N_{i}^{-}$ are the sets of positive and negative neighbors of $i$ respectively. Inspired by [Bresson et al., 2024], we integrate the Kolmogorov-Arnold layer (KAN) in the signed graph convolutional neural network (SGCN) by replacing the activation function $\sigma$ and the weight matrix $W$ with the KAN layer in the formulation. This improvement is easily integrated as the KAN layers have trainable parameters and introduce non-linearity, superseding the roles of the activation function and the weight matrix. Therefore, the Kolmogorov-Arnold signed graph convolutional network (KASGCN) can be formulated as follows:
@@
$h_{i}^{B(1)}=\phi^{B(1)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (9)
$h_{i}^{U(1)}=\phi^{U(1)}([\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (10)
@@
$h_{i}^{B(l)}=\phi^{B(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}])$ (11)
$h_{i}^{U(l)}=\phi^{U(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}])$ (12)
where $\phi^{B(l)}$ and $\phi^{U(l)}$ here represent the KAN layer used to update the positive and negative representations, respectively. This layer can be any state-of-the-art variant such as OriginalKAN $\phi_{o}$ [Liu et al., 2024], FourierKAN $\phi_{F}$ [Xu et al., 2024], LaplaceKAN $\phi_{l}$ [Yin et al., 2024], and WaveletKAN $\phi_{w}$ [Bozorgasl and Chen, 2024]. We later experimented with these variants to determine how each performed when integrated into SGCN.
4 Empirical Evaluation

### v3

@@ Abstract
Graph Representation Learning aims to create effective embeddings for nodes and edges that encapsulate their features and relationships. Graph Neural Networks (GNNs) leverage neural networks to model complex graph structures. Recently, the Kolmogorov-Arnold Neural Network (KAN) has emerged as a promising alternative to the traditional Multilayer Perceptron (MLP), offering improved accuracy and interpretability with fewer parameters. In this paper, we propose the integration of KANs into Signed Graph Convolutional Networks (SGCNs), leading to the development of KAN-enhanced SGCNs (KASGCN). We evaluate KASGCN on tasks such as signed community detection and link sign prediction to improve embedding quality in signed networks. Our experimental results indicate that KASGCN exhibits competitive or comparable performance to standard SGCNs across the tasks evaluated, with performance variability depending on the specific characteristics of the signed graph and the choice of parameter settings. These findings suggest that KASGCNs hold promise for enhancing signed graph analysis with context-dependent effectiveness.
@@
Graph Representation Learning is creating embeddings for nodes and edges based on the features and connections in a graph. These embeddings capture the graph’s local and global features and are useful for node classification and link prediction tasks. One of the most well-known methods for learning graph embeddings is through Graph Neural Networks (GNNs) [Xu, 2021]. GNNs have proven to understand complex graph relationships and structures effectively. Graph neural networks gather information from neighboring nodes, gradually refining the embeddings through multiple layers [Fan et al., 2019, Bessadok et al., 2023]. The Kolmogorov-Arnold Neural Network (KAN) has emerged recently as an alternative to the traditional Multilayer Perceptron (MLP) architecture [Liu et al., 2024]. In contrast to traditional multilayer perceptrons (MLPs), KAN utilizes learnable univariate functions instead of fixed activation functions. KANs have outperformed MLPs for the numerical analysis and partial differential equation solving [Ji et al., 2024] as the learnable activation functions on edges show more flexibility in learning complex models, and KAN can also handle high-dimensional data. The fusion of KANs and graph neural networks in several ways has improved the performance and the quality of the GNN embeddings [De Carlo et al., 2024, Kiamari et al., 2024, Bresson et al., 2024] in the node classification, link prediction, and graph classification tasks.
This paper introduces the first known integration of the Kolmogorov-Arnold Neural Network (KAN) and its variants into signed Graph Neural Networks (GNNs), evaluating their impact on tasks such as signed clustering and link sign prediction. Additionally, we investigate how KANs can enhance the quality of embeddings generated by signed GNNs. Signed GNNs are particularly suited for graphs where edges represent positive or negative interactions, capturing the dynamics of various relationships. The Signed Graph Convolutional Networks (SGCNs) method, proposed by Derr et al., adapts Graph Convolutional Networks to signed networks by leveraging balance theory [Derr et al., 2018]. Balance theory, which addresses the dynamics of attitudes within networks, has applications in edge sentiment prediction and anomaly detection. In this context, a signed network is considered strongly balanced if every fundamental cycle contains an even number of negative edges.
@@
Signed Graph Convolutional Neural Network (SGCN) is the baseline and the most robust variant of signed GNNs in the field, as it adapts the Graph Convolutional Neural Network for signed networks based on the balance theory [Derr et al., 2018]. Balance theory is pivotal in explaining attitudes’ evolution within signed graph networks. Heider formalized the balance theory in [Abelson and Rosenberg, 1958], and Harary introduced the mathematical formulation and the $k$-way balancing [Cartwright and Harary, 1956, Harary and Cartwright, 1968]. The balance theory has been widely applied in various domains, including edge sentiment prediction, content and product recommendations, and anomaly detection [Derr et al., 2020, Garimella et al., 2021, Interian et al., 2022, Amelkin and Singh, 2019]. In the context of signed networks, a network is considered strongly balanced if every fundamental cycle consists of an even number of negative edges. The Signed Graph Convolutional Network (SGCN) incorporates the balance theory by maintaining two distinct representations at each layer: one for the suggested friends, where the path connecting the node contains an even number of negative links, and another for the suggested enemies, where the path includes an odd number of negative links.
Figure 1: Illustration of SGCN and KAN-enhanced SGCN of a signed network of 4 nodes and three edges.
@@
This research project is the first to integrate KAN (with its variants) into signed graph neural networks and empirically evaluate its impact on performance on two downstream tasks: signed clustering and link sign prediction for signed networks. Next, the project outlines KANs’ potential to increase the expressive power of the embeddings generated by signed graph neural networks. Section 0.1 discusses the related work regarding KAN integration into several architectures, their uses, and the background on KAN and SGCN. In Section 1.3, we present the methodology of our work and pose research questions for this study. In Section 3, we empirically evaluate the KAN-integrated signed GNN and its variants in several tasks, including efficiency examination and embedding comparison. We structure this section by presenting the hypothesis, results, and observations. In Section 4, we summarize our findings.
0.1 Related Work
@@
Cheon et al. proposed a method that combines Kolmogorov-Arnold Networks (KANs) with pre-trained Convolutional Neural Networks (CNNs), specifically VGG16 and MobileNetV2 models, for scene classification in satellite imagery [Cheon, 2024]. The proposed approach demonstrated high accuracy, requiring fewer epochs and parameters than traditional Multilayer Perceptrons (MLPs) that do not integrate KANs [Cheon, 2024]. KANs have been explored for data representation using autoencoders and compared against conventional Convolutional Neural Networks (CNNs) on datasets such as MNIST, SVHN, and CIFAR-10. And it is shown that KAN-based autoencoders deliver competitive reconstruction accuracy [Moradi et al., 2024]. Dong et al. evaluated the performance of KANs on time series data and found that KANs can match or even surpass the performance of MLP across 128 different time series datasets [Dong et al., 2024]. Lu et al. assessed using KANs for fraud detection and concluded that their performance varies depending on the context [Lu and Zhan, 2024]. The study in [Anonymous, 2024] presented a proof of concept for employing KANs in graphs to predict molecules’ binding affinity to protein targets. The model in this study did not attain state-of-the-art performance, but it emphasized its promising method in computational drug discovery. For the Graph Neural Network integration, Bresson et al. integrated KAN layers into unsigned graph neural networks (GNNs) for the node and graph classification tasks [Bresson et al., 2024], and Li et al. integrated Fourier series-based KANs to optimize GNNs further for the molecular property prediction. [Li et al., 2024].
1 Introducing the Kolmogorov Arnold Signed Graph Convolutional Networks
1.1 Kolmogorov-Arnold Networks
@@
The Kolmogorov-Arnold representation theorem states that any multivariate function can be formulated as a combination of continuous univariate functions [Kolmogorov, 1956], as defined in Eq. 1.
@@
The parameters $\alpha$ and $\phi$ are the univariate functions and $d$ is the dimension of the input. The output is an aggregated output of each input after it passes through an univariate and non-linear spline function. The main hyperparameters for the spline function are the spline order (degree of B-splines) and the grid size (number of intervals to approximate the real function).
Kolmogorov-Arnold Networks (KANs) demonstrate greater expressiveness than standard multilayer perceptrons (MLPs) and superior performance with significantly fewer parameters [Liu et al., 2024]. Where MLPs adjust the weights globally based on the training data, KANs spline control points impact only local regions, thereby preserving more information [Fong and Seidel, 1991]. Next, $\phi(x)$ from Eq. 1 is defined in Eq. 2.
$\phi(x)=w_{b}*\text{SiLU}(x)+w_{s}*\text{spline}(x).$ (2)
The $\text{SiLU}(x)$ is the SiLU activation function $x/(1+e)$ and $\text{spline}(x)$ is defined as the is a linear combination of B-splines, $\text{spline}(x)=\sum_{i}c_{i}B_{i}(x)$ where $c_{i},w_{b}$, and $w_{s}$ are trainable and control the magnitude of the activation function. $w_{s}=1$ and $w_{b}$ are initialized using Xavier initialization.
1.2 Signed Graph Convolutional Neural Networks
Derr et al. apply balance theory to effectively aggregate features from a node’s neighbors in signed networks [Derr et al., 2018]. Each node has a positive embedding and a negative embedding representation. During the message-passing process, the positive embedding updates by aggregating the positive embeddings from its positive neighbors. The negative embeddings update from its negative neighbors, concatenating the resulting aggregate to the node’s positive representation. By aggregating the negative embeddings from positive neighbors and the positive embeddings from negative neighbors concatenate with the node’s negative representation.
Conversely, the negative embedding is updated. The final node embedding is obtained by concatenating both the positive and negative embeddings. The signed graph convolutional network (GCN) formulation is described as follows.
@@ 1.2 Signed Graph Convolutional Neural Networks
$h_{i}^{B(1)}=\sigma(W^{B(1)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$
$h_{i}^{U(1)}=\sigma(W^{U(1)}[\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$
@@
$h_{i}^{B(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}])$
$h_{i}^{U(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}])$
The number of layers is $l$, $h_{i}^{B(l)}$ and $h_{i}^{U(l)}$ are the positive and negative representations of node $i$ in layer $l$ respectively, $\sigma$ is an activation function, $W^{B(l)}$ and $W^{U(l)}$ are the weight matrices for updating the positive and negative representations respectively, $N_{i}^{+}$ and $N_{i}^{-}$ are the sets of positive and negative neighbors of $i$ respectively.
1.3 Outline of the KASGCN Method
In this section, we introduce the Kolmogorov-Arnold Signed Graph Convolutional Network (KASGCN) method to integrate the Kolmogorov-Arnold layer (KAN) in the signed graph convolutional neural network (SGCN) by replacing the weight matrix $W$ with the KAN layer in the formulation. The KAN layers have trainable parameters and introduce non-linearity, superseding the role of the weight matrix illustrated for the unsigned example in [Bresson et al., 2024]. We formalize the Kolmogorov-Arnold signed graph convolutional network (KASGCN) method as follows:
@@ 1.3 Outline of the KASGCN Method
$h_{i}^{B(1)}=\sigma(\phi^{B(1)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}]))$
$h_{i}^{U(1)}=\sigma(\phi^{U(1)}([\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}]))$
@@
$h_{i}^{B(l)}=\sigma(\phi^{B(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}]))$
$h_{i}^{U(l)}=\sigma(\phi^{U(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}]))$
the $\phi^{B(l)}$ and $\phi^{U(l)}$ here represent the KAN layer used to update the positive and negative representations, respectively. This layer can be any state-of-the-art variant such as OriginalKAN $\phi_{o}$ [Liu et al., 2024], FourierKAN $\phi_{f}$ [Xu et al., 2024], LaplaceKAN $\phi_{l}$ [Yin et al., 2024], and WaveletKAN $\phi_{w}$ [Bozorgasl and Chen, 2024]. We later experimented with these variants to determine how each performed when integrated into SGCN. As an illustration, Figure 1 depicts both the SGCN and the KAN-enhanced SGCN starting from $l=1$. In addition, we pose the following research questions, and we answer them in Section 3:
  1. 1.
    Does incorporating the KAN layer into the signed graph convolutional network improve the clustering quality when using Kmeans++ on the embeddings?
  2. 2.
    Do the embeddings generated from integrating the KAN layer into SGCN increase the predictive power of logistic regression for sign classification?
  3. 3.
    How efficient is KASGCN compared to SGCN?
  4. 4.
    How similar do SGCN and KASGCN produce the embeddings?
  5. 5.
    Does integrating other KAN variants into SGCN perform just as well as integrating the original KAN layer?
2 Proof Of Concept Setup